the number of gram molecules of oxygen in 6.023*10^24CO molecules is -

Grade:11

3 Answers

Harshit
6 years ago
Number of gram molecule of Oxygen =Number of mole of Oxygen gas . Moles of CO=6.023 × 10^24/6.023 × 10^23 = 10. Number of mole oxygen atom=10 . No of gram molecule of oxygen gas = 10/2 =5
RAJPAL DEEPAK MORE
5 years ago
Given no of molecule of CO is 6.02×10^24.moles of CO is =6.02×10^24/6.02×10^23=10. Moles of O2 atom is 10 so moles of O atom is 10/2=5g.
 
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Number of gram molecule of Oxygen =Number of mole of Oxygen gas .
Moles of CO = 6.023 × 10^24/6.023 × 10^23 = 10.
Number of mole oxygen atom=10 .
No of gram molecule of oxygen gas = 10/2 =5

Thanks and Regards